# Quantum-enhanced mean value estimation via adaptive measurement

### v3

@@ 5 Conclusions
Acknowledgements: K.W. and K.F. thank IPA for the support through MITOU target program. We would like to thank Dr. Yasunari Suzuki, Dr. Yuuki Tokunaga, and Dr. Tomoki Tanaka for helpful discussions. This work was supported by MEXT Quantum Leap Flagship Program Grant Number JPMXS0118067285 and JPMXS0120319794.
@@ Appendix E Proof for statistical properties
$=\sum_{\alpha^{\prime}_{m}\in D_{m-1}}\sum_{\begin{subarray}{c}\bm{x}_{m-1}:\\ \alpha_{m}(\bm{x}_{m-1})=\alpha^{\prime}_{m}\end{subarray}}\mathcal{L}_{m-1}\sum_{{x}^{(m)}}F_{m}\left(x^{(m)};\alpha^{\prime}_{m},\theta^{*}\right)\mathbf{1}{\left\{F_{m}\left(X^{(m)};\alpha_{m}^{*},\theta^{*}\right)>F_{m}\left(X^{(m)};\alpha_{m}^{*},\theta\right)\right\}}({x}^{(m)})$ (E.17)
@@
where $\mathcal{N}$ denotes a centered normal distribution with variance $\sin^{2}{\theta^{*}}$ and normal-$\to$ means the convergence in distribution as $N$ increases.
@@
$=\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})=\alpha_{k}^{*}~{}\forall k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$ (E.44)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$ (E.47)
$=\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})=\alpha_{k}^{*}~{}\forall k\end{subarray}}\prod_{m=1}^{M}F_{m}\left(x^{(m)};\alpha_{m}^{*},\theta^{*}\right)h(\bm{x}_{M},\alpha^{*}_{2},\cdots,\alpha^{*}_{M})$ (E.50)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$ (E.53)
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$ (E.56)
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})\left|h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))\right|$ (E.62)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\prod_{m=1}^{M}F_{m}\left(x^{(m)};\alpha_{m}^{*},\theta^{*}\right)\left|h(\bm{x}_{M},\alpha^{*}_{2},\cdots,\alpha^{*}_{M})\right|$ (E.65)

### v4

@@ 5 Conclusions
Acknowledgements: K.W. and K.F. thank IPA for the support through MITOU target program. We would like to thank Dr. Yasunari Suzuki, Dr. Yuuki Tokunaga, and Dr. Tomoki Tanaka for helpful discussions. K.W. was supported by JST SPRING, Grant Number JPMJSP2123. This work was supported by MEXT Quantum Leap Flagship Program Grant Number JPMXS0118067285 and JPMXS0120319794.
@@ Appendix E Proof for statistical properties
$=\sum_{\alpha^{\prime}_{m}\in D_{m-1}}\sum_{\begin{subarray}{c}\bm{x}_{m-1}:\\ \alpha_{m}(\bm{x}_{m-1})=\alpha^{\prime}_{m}\end{subarray}}\mathcal{L}_{m-1}\sum_{{x}^{(m)}}F_{m}\left(x^{(m)};\alpha^{\prime}_{m},\theta^{*}\right)\mathbf{1}{\left\{F_{m}\left(X^{(m)};\alpha_{m}^{*},\theta^{*}\right)>F_{m}\left(X^{(m)};\alpha_{m}^{*},\theta\right)\right\}}({x}^{(m)})$
@@
where $\mathcal{N}$ denotes a centered normal distribution with variance $\sin^{2}{\theta^{*}}$ and $\to$ means the convergence in distribution as $N$ increases.
@@
$=\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})=\alpha_{k}^{*}~{}\forall k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$
$=\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})=\alpha_{k}^{*}~{}\forall k\end{subarray}}\prod_{m=1}^{M}F_{m}\left(x^{(m)};\alpha_{m}^{*},\theta^{*}\right)h(\bm{x}_{M},\alpha^{*}_{2},\cdots,\alpha^{*}_{M})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\mathcal{L}_{M}(\bm{x}_{M};\theta^{*})\left|h(\bm{x}_{M},\alpha_{2}(\bm{x}_{1}),\cdots,\alpha_{M}(\bm{x}_{M-1}))\right|$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{\begin{subarray}{c}\bm{x}_{M}\\ {\alpha}_{k}(\bm{x}_{k-1})\neq\alpha_{k}^{*}~{}\exists k\end{subarray}}\prod_{m=1}^{M}F_{m}\left(x^{(m)};\alpha_{m}^{*},\theta^{*}\right)\left|h(\bm{x}_{M},\alpha^{*}_{2},\cdots,\alpha^{*}_{M})\right|$